OpenStudy (experimentx):

Solve: \[ x^2 (a - bx) {dy^2 \over dx^2} - x (5a - 4bx) {dy \over dx} + 3(2a - bx)y = 6a^2\] Solution (according to answer sheet): \[ y(a-bx) = Ax^2 + Bx^3 + C\] Looking for right substitution to convert it into nice form: \[ u^2 {dv^2 \over du^2} + P_1 u {dv \over du} + P_2 v = F(u)\]

OpenStudy (experimentx):

also more thread on this Q http://math.stackexchange.com/questions/176220/need-hints-to-solve-this-de

OpenStudy (anonymous):

for the last equation ... suppose that \(u=x\) and \(v=y(a-bx)\) , \(P_1=-4\) , \(P_2=6\) and \(F(u)=constant\)

OpenStudy (anonymous):

then ur equation will be \[x^2 \frac{d^2v}{dx^2}-4x \frac{dv}{dx}+6v=C \] Using Euler method answer for this equation would be \[v=Ax^2+Bx^3+C \]

OpenStudy (anonymous):

i tried to sub \(v=y(a-bx)\) in the original equation but it does'nt work....:(

OpenStudy (experimentx):

let me verify the solution using Maple or Mathematica ...

OpenStudy (anonymous):

ok...:)

OpenStudy (anonymous):

Let \( a=b=1\) in the ODE, then \[ (1-x) x^2 y''(x)-(5-4 x) x y'(x)+3 (2-x) y(x)=6 \] If we take \[ y(x) =\frac{A x^2+B x^3+C}{1-x} \] and we replace in the original ODE and multiply the two sides by \((1-x)^2\), we get by Mathematica \[x^4 (7 B-A)+x^3 (4 A-3 B)+x^2 (9 C-2 A)-3 B x^5-14 C x+6 C=6 (x-1)^2\] We get to solve \[ 4 A-3 B=0\\ -3 B=0\\ 7B-A=0\\ 12-14C=0\\ 6 C -6=0\\ -2 A+9C -C =0 \] There is no solution for the system above.

OpenStudy (anonymous):

yeah i think there is something wrong with that...

OpenStudy (experimentx):

i guess so ... the problem in book might be wrong. I'm also getting strange answer using Maple.